The Napa Co. just paid a dividend of $3.15 per share on its stock. The dividends are expected to grow at a constant rate of 5 percent per year indefinitely. If investors require a return of 12 percent on the stock, what is the current price? What will the price be in four years?
Answer #1

The formula to calculate current price is given below:

Current\, price=\frac{Last\, dividend\times \left ( 1+Growth \right )}{Required\, return-Growth}

Current\, price=\frac{3.15\times \left ( 1+0.05 \right )}{0.12-0.05}

Current\, price=\frac{3.3075}{0.07}

Current\, price=47.25

Current price is $47.25.

The dividend in 5 years will be:

Dividend=3.15\times \left ( 1+0.05 \right )^{5}

Dividend=3.15\times 1.276281563

Dividend=4.020286922

Price in 4 years will be:

Share\, price=\frac{4.020286922}{0.12-0.05}

Share\, price=57.43